Flagstaff Enterprises expected to have free cash flow in the coming year of $8 million, and this free cash flow is expected to grow at a rate of 3% per year thereafter. Flagstaff has an equity cost of capital of 13%, a debt cost of capital of 7%, and it is in the 35% corporate tax bracket. If Flagstaff currently maintains a 0.5 debt to equity ratio, then the value of Flagstaffʹs interest tax shield is closest to (all calculations round up to two decimals places) :
Answer #1

RWACC = rE + rD (pre-tax)

RWACC = 0.13+ 0.07= 0.1033

VU = $109.09 million

RWACC = rE + rD (1-tc) after tax

RWACC = 0.13 + 0. 07 (1-0.35) = 0.09244

VL= $128.11 million

present tax shield = Value of levered-Value of unlevered = $128.11- $109.09 = $19.02